What is the new freezing point of a 0.50 m solution of a nonelectrolyte dissolved in water?


1
Expert's answer
2021-04-07T11:31:38-0400

Solution:

The freezing point depression (ΔTf) is related to the molality (m) and the freezing point depression constant (Kf) by ΔTf = K× m (for nonelectrolyte).


Kf(water)=1.86˚C/m

m = 0.50 m

Hence,

ΔTf = (1.86˚C/m) × (0.50 m) = 0.93˚C

Water normally freezes at 0˚C.

Hence,

Freezing point = 0˚C - 0.93˚C = -0.93˚C

Freezing point = -0.93˚C


Answer: The new freezing point of a solution is -0.93˚C.